A sample contains H atom in first excited state, He+ in third excited state, Li2+ in fifth excited state. Find the number of spectral lines produced when all electrons return to ground state . ..... ....    .. ......... ANS:20

Grade:12th pass

1 Answers

Shweta Ahire
5 years ago
When electrons come to ground state then 
No of spectral lines = n( n-1)/2.....n is no. Of orbit in which electron is present initially 
 
So for H atom it is 2(2-1)/2=1 
For He+ , 4(4-1)/2 = 6
For Li2+,  6(6-1)/2=15
So total spectral lines are 22